# Equivariant concordance of periodic 2-knots in $S^{4}$

### v1

@@ 1.2. Equivariant concordance group structure
A direction for a $d$-periodic 2-knot $(S,\tau)$ is a choice of orientation on $\operatorname{Fix}(\tau)$ together with a choice of labeling of the two points in $\operatorname{Fix}(\tau)\cap S$ with the labels $p_{+}$ and $p_{-}$.
@@
Let $(S,\tau)$ and $(S^{\prime},\tau^{\prime})$ be two directed $d$-periodic 2-knots. Then we define their equivariant connected sum, denoted $(S\#S^{\prime},\tau\#\tau^{\prime})$ or simply $S\#_{\mathbb{Z}/d\mathbb{Z}}S^{\prime}$, to be the $d$-periodic 2-knot obtained as follows: First, remove an equivariant neighborhood of $p_{+}$ in $S^{4}$ from the copy of $S^{4}$ containing $S$, and an equivariant neighborhood of $p_{-}^{\prime}$ from the copy of $S^{4}$ containing $S^{\prime}$. Then glue these two copies of $S^{4}$ together along their boundaries such that:
@@
    the intersections of $\operatorname{Fix}(\tau)$ and $\operatorname{Fix}(\tau^{\prime})$ with the deleted neighborhoods are glued together along their common boundary with opposite orientations (so that the union of $\operatorname{Fix}(\tau)$ with $\operatorname{Fix}(\tau^{\prime})$ has a consistent orientation given by the chosen orientations on each of its components),
@@
    $\tau$ and $\tau^{\prime}$ agree on the common boundary of the punctured 4-spheres, and
@@
    the punctured copies of $S$ and $S^{\prime}$ are glued together along their boundaries.
The resulting 2-knot $S\#_{\mathbb{Z}/d\mathbb{Z}}S^{\prime}$ will then be $d$-periodic under the action given by the actions $\tau$ and $\tau^{\prime}$ on their constituent components.
As in the case with strongly invertible classical knots, equivariant connected sum is not usually commutative; that is, $S\#_{\mathbb{Z}/d\mathbb{Z}}S^{\prime}$ may not be equivariantly isotopic to $S^{\prime}\#_{\mathbb{Z}/d\mathbb{Z}}S$, where we fix an equivariant map identifying their ambient copies of $S^{4}$ that preserves the orientation on the fixed set. (In the classical dimension, Di Prisa [DP23] has shown that the equivariant concordance group of strongly invertible knots is not itself commutative, and so equivariant connected sum is not even commutative up to concordance. However, we will see that this is not the case for $\mathcal{C}^{2}_{d}$.)
For the following definition, extend the linear order $d$ rotation $\rho$ on $S^{2}$ to $S^{2}\times I$ to $\hat{\rho}$ such that $\hat{\rho}(x,t)=(\rho(x),t)$.
@@
$\tau\circ\Psi=\Psi\circ\hat{\rho}$
We will work primarily with $B^{5}$ and think of the concordance taking place inside an embedded copy of $S^{4}\times I$ such that the actions on each $S^{4}$ are the restriction of a global order $d$ rotation on $B^{5}$ fixing a properly embedded unknotted 3-ball. If $(S,\tau)$ and $(S^{\prime},\tau^{\prime})$ are directed, we say that a concordance between them is directed if the orientations on $\operatorname{Fix}(\tau)$ and $\operatorname{Fix}(\tau^{\prime})$ agree. Directed equivariant concordance is then an equivalence relation on the set of directed periodic 2-knots.
@@
Let $C^{2}_{d}$ be the set of directed $d$-periodic 2-knots. The equivariant concordance group $\mathcal{C}^{2}_{d}$ of $d$-periodic 2-knots is the quotient of the set $C^{2}_{d}$ by directed equivariant concordance, with group multiplication given by equivariant connected sum.

### v2

@@ 1.2. Montesinos annuli
A smooth Montesinos annulus is a copy of $S^{1}\times I$ smoothly and properly embedded in $S^{1}\times B^{3}$ such that $\partial N\cong S^{1}\sqcup S^{1}$ are mapped to disjoint copies of $S^{1}\times\{\text{pt}\}\subset S^{1}\times S^{2}\cong\partial(S^{1}\times B^{3})$.
We choose this name for these annuli in light of the following setup: Consider $R$ a smoothly unknotted copy of $S^{2}$ inside of $S^{4}$, and let $S$ be a smooth 2-knot in $S^{4}$ intersecting $R$ transversely such that $S\cap R$ consists of exactly two points. The pair $(R,S)$ is often called a Montesinos twin, albeit one where one of the twins is unknotted. Now consider removing a tubular neighborhood $\nu(R)$ from $S^{4}$, including neighborhoods of its intersection points with $S$. What is left is an annulus embedded in $S^{4}\setminus(D^{2}\times S^{2})\cong S^{1}\times B^{3}$ exactly as described in Definition 1.4.
With this in mind, we associate to each Montesinos annulus $N$ a corresponding sphere $S$ and a removed sphere $R$, such that $N\subset S^{1}\times B^{3}$ is obtained by removing $\nu(R)$ from $S\subset S^{4}$. It is important for this construction that we fix coordinates on the copy of $S^{1}\times B^{3}$ in which the Montesinos annulus lives (and thus an identification of the boundary with $S^{1}\times S^{2}$), so that the corresponding sphere is uniquely determined. Otherwise, the Gluck map introduces an ambiguity when gluing back in the removed sphere. Throughout the paper, we implicitly fix such an identification.
We say that a Montesinos annlulus is unknotted if it is boundary parallel in $S^{1}\times S^{3}$. Note that the fundamental group of the complement of the unknotted annulus is $\mathbb{Z}^{2}$.
[…]
Figure 1. Two projections of an unknotted Montesinos annulus and its removed sphere. On the left, we see the annulus $N$ with its boundary components on $\partial(\nu(R))$. On the right and in a different 3-dimensional projection, we see $R$ locally as a plane, and one slice of the corresponding sphere $S$ for NN as a circle intersecting $R$ in two points.
Example 1.5.
Let $S_{T}$ be the one-twist spun trefoil 2-knot inside $S^{4}$ (not up to isotopy), and let $R$ be an unknotted copy of $S^{2}$ representing its spin axis. Remove a neighborhood $\nu(R)$ from $S^{4}$, including neighborhoods of its intersection with $S_{T}$. This leaves us with a Montesinos annulus inside of $S^{1}\times B^{3}$, which we will call $N_{T}$.
The key point to illustrate is that although the sphere $S_{T}$ is isotopic to the unknot in $S^{4}$ (see Zeeman [ZEE65]), the annulus $N_{T}$ is not isotopic to the unknotted annulus rel. boundary. It is distinguished by the fundamental group of its complement in $S^{1}\times B^{3}$, which is isomorphic to the direct product of the trefoil knot group with $\mathbb{Z}$.
Note also the following construction, illustrated in Figure 2, which will be important later. Following Zeeman, we can construct a 3-ball $B$ in $S^{4}$ bounded by $S_{T}$ by twist-spinning a relative Seifert surface $F$ for the trefoil tangle $T$. Close inspection of this construction shows that the isotopy type of the spin axis $A$ inside of this 3-ball is itself also a trefoil. The Seifert surfaces for $T$ together form a genus one fibration of the complement of $A$ in $B$ which glue together with precisely the trefoil monodromy, as shown in Figure 2. Hence the intersection of $B$ with the removed sphere $R$ for $N_{T}$ is also a trefoil tangle.
[…]
Figure 2. A movie of the one-twist spun trefoil, illustrated by twisting an the spin axis AA on the boundary while fixing the trefoil tangle $T$. The Seifert surfaces shown patch together to form a 3-ball in which the complement of $A$ is fibered by punctured tori with the trefoil monodromy.
Definition 1.6.
Two Montesinos annuli $N$ and $N^{\prime}$ are smoothly concordant if there exists $E$ a copy of $S^{1}\times I\times I$ smoothly and properly embedded into $S^{1}\times B^{3}\times I$ such that $E|_{S^{1}\times I\times\{0\}}=N$, $E|_{S^{1}\times I\times\{1\}}=N^{\prime}$, and $E|_{S^{1}\times\{0,1\}\times I}=S^{1}\times\{p_{+},p_{-}\}\times I$ for some fixed points $p_{+},p_{-}$ in $S^{1}\times S^{2}$.
Equivalently, $N$ and $N^{\prime}$ are concordant if their corresponding spheres $S$ and $S^{\prime}$ are concordant in $S^{4}\times I$ such that the intersection of the concordance with $R\times I\subset S^{4}\times I$ consists of two arcs $\{p_{+}\}\times I,\{p_{-}\}\times I$ which are embedded in $R\times I$ by a level preserving map. We say that an annulus $N$ is slice if it is concordant to the unknotted annulus.
Remark 1.7.
Note that the kind of relative concordance for annuli we consider requires that both boundary components of the annulus stay fixed throughout the whole concordance, and are not allowed to vary in their isotopy class. Because of this restriction, we may obtain a different characterization of slice annuli adapted to the setting of $B^{5}$. Parametrize a neighborhood of $B^{5}$ near its boundary with coordinates $(x,y,z,w,t)$ such that $t\geq 0$. Consider $S^{1}\times B^{3}\times I$ sitting inside of $B^{5}$ such that $S^{1}\times B^{3}\times\{0\}$ is contained in $\partial B^{5}$, and so that projection onto the $I$ factor agrees with projection onto the $t$ coordinate. If $N$ is slice, then we may consider it as contained in $S^{1}\times B^{3}\times\{0\}$, and the unknotted annulus $U$ as contained in $S^{1}\times B^{3}\times\{1\}$. By filling in a 3-ball bounded by the corresponding sphere for $U$ lying in the $t=1$ slice (which necessarily intersects the removed sphere $R$ in that slice in an unknotted arc) and perturbing the 3-ball slightly (in a way formalized in Lemma 5.3), we see that the corresponding sphere $S$ for $N$ bounds a 3-ball into $B^{5}$ whose intersection with an unknotted copy $Q$ of $B^{3}$ has exactly one critical point with respect to projection onto the $t$ coordinate. We call such a slice ball minus a neighborhood of $Q$ a slice cylinder for $N$. This characterization is the one we will work with most often, and many of our constructions and proofs take place in this parametrization.
Throughout this paper, given a Montesinos annulus $N$, $R$ will always stand for its removed sphere, and $S$ for its corresponding sphere. For brevity, we will also often drop the word “Montesinos” from Montesinos annulus, where it is clear from context.
1.3. Concordance group structure
1.3.1. Concordance of annuli
To define a group structure on the set of Montesinos annuli up to concordance, we require a notion of connected sum, which in turn necessitates some extra data. We take inspiration from the connected sum for strongly invertible classical knots first defined by Sakuma in [SAK86] in our definition. Throughout this subsection, $N$ is a Montesinos annulus, and $S$ and $R$ are its corresponding and removed spheres, respectively.
Definition 1.8.
@@ 1.3.1. Concordance of annuli
A direction for an annulus $N$ is a choice of orientation on $R$ together with a choice of labeling of the two boundary components of $N$ with the labels $c_{+}$ and $c_{-}$.
We call an annulus with a choice of direction a directed annulus. The concordance of annuli given in Definition 1.6 can then be refined to a notion of directed concordance, in which we require that the directions of the two annuli on either end of the concordance agree.
Definition 1.9.
Let $N$ and $N^{\prime}$ be two directed annuli, and let $R$ and $R^{\prime}$ denote the removed spheres in their respective copies of $S^{4}$. Then we define their stack, denoted $N\natural N^{\prime}$, to be the annulus obtained as follows: From the copy of $S^{4}$ containing $N$, remove a ball such that $c_{+}$ lies in the new $S^{3}$ boundary component. From the copy of $S^{4}$ containing $N^{\prime}$, remove a ball such that $c_{-}^{\prime}$ lies in the new $S^{3}$ boundary component. Then glue these two copies of punctured $S^{4}$ together along their boundaries such that:
  1. (1)
    the intersections of $R$ and $R^{\prime}$ with the deleted neighborhoods are glued together along their common boundary with opposite orientations, so that the resulting sphere $R\#R^{\prime}$ has a consistent orientation, and
  2. (2)
    NN and $N^{\prime}$ are glued so that $c_{+}$ is identified with $c_{-}^{\prime}$ with opposite orientations.
The result is an annulus $N\natural N^{\prime}$ whose corresponding sphere is the connected sum $S\#S^{\prime}$ which intersects the connected sum of removed spheres $R\#R^{\prime}$ in two points.
[…]
Figure 3. A schematic of one annulus stacked on top of another.
As in the case with strongly invertible classical knots, annulus connected sum is not usually commutative; that is, $N\#N^{\prime}$ may not be isotopic to $N^{\prime}\#N$. In comparison with the classical dimension, note Di Prisa [DI 23] has shown that the equivariant concordance group of strongly invertible knots is not itself commutative, and so equivariant connected sum is not even commutative up to concordance. However, we will see that this is not the case for the concordance group for Montesinos annuli $\mathcal{C}_{2}^{\circ}$.
Definition 1.10.
Let $C_{2}^{\circ}$ be the set of directed Montesinos annuli. The relative concordance group $\mathcal{C}_{2}^{\circ}$ of Montesinos annuli is the quotient of the set $C_{2}^{\circ}$ by directed concordance, with group multiplication given by stacking.
1.3.2. Equivariant concordance
The situation for periodic 2-knots is almost identical. We reuse much of the same language.
Definition 1.11.
A direction for a $d$-periodic 2-knot $(\tilde{S},\rho)$ is a choice of orientation on $\operatorname{Fix}(\rho)$ together with a choice of labeling of the two points in $\operatorname{Fix}(\rho)\cap\tilde{S}$ with the labels $\hat{p}_{+}$ and $\hat{p}_{-}$.
@@ 1.3.2. Equivariant concordance
Let $(\tilde{S},\rho)$ and $(\tilde{S}^{\prime},\rho^{\prime})$ be two directed $d$-periodic 2-knots. Then we define their equivariant connected sum, denoted $(\tilde{S}\#\tilde{S}^{\prime},\rho\#\rho^{\prime})$ or just $\tilde{S}\#_{\mathbb{Z}/d\mathbb{Z}}\tilde{S}^{\prime}$, to be the $d$-periodic 2-knot obtained as follows: First, remove an equivariant neighborhood of $p_{+}$ in $S^{4}$ from the copy of $S^{4}$ containing $\tilde{S}$, and an equivariant neighborhood of $p_{-}^{\prime}$ from the copy of $S^{4}$ containing $\tilde{S}^{\prime}$. Then glue these two copies of $S^{4}$ together along their boundaries such that:
@@
    the intersections of $\operatorname{Fix}(\rho)$ and $\operatorname{Fix}(\rho^{\prime})$ with the deleted neighborhoods are glued together along their common boundary with opposite orientations (so that the union of $\operatorname{Fix}(\rho)$ with $\operatorname{Fix}(\rho^{\prime})$ has a consistent orientation given by the chosen orientations on each of its components),
@@
    $\rho$ and $\rho^{\prime}$ agree on the common boundary of the punctured 4-spheres, and
@@
    the punctured copies of $\tilde{S}$ and $\tilde{S}^{\prime}$ are glued together along their boundaries.
The resulting 2-knot $\tilde{S}\#_{\mathbb{Z}/d\mathbb{Z}}\tilde{S}^{\prime}$ will then be $d$-periodic under the action given by the actions $\rho$ and $\rho^{\prime}$ on their constituent components.
For the following definition, extend the linear order $d$ rotation $\sigma$ on $S^{2}$ to $S^{2}\times I$ to the action $\tau$ such that $\tau(x,t)=(\sigma(x),t)$.
@@
$\rho\circ\Psi=\Psi\circ\tau$
satisfying the condition that
$\Psi|_{\operatorname{Fix}(\tau)}(x,t)=(\psi(x),t)$
for some map $\psi\colon\operatorname{Fix}(\sigma)\to\operatorname{Fix}(\rho)$.
Remark 1.14.
This second condition amounts to requiring that our equivariant concordances be level preserving on the fixed sets of the action. This is equivalent to the boundary conditions in Definition 1.4 for concordances of Montesinos annuli, and is necessary in order for the equivariant connected sum operation defined above to act as a group multiplication for equivariant concordance classes. In principle, one could study equivariant concordances or equivariant sliceness without this condition, but it is not clear that the set of equivalence classes under a looser definition of equivariant concordance actually forms a group. Also, removing this requirement complicates many Morse-theoretic arguments which are central to the results in this paper, and it is not immediately obvious how to recover these results without them.
We will work primarily with $\tilde{B}^{5}$ and think of the equivariant concordance as taking place inside an embedded copy of $\tilde{S}^{4}\times I$ such that the actions on each $\tilde{S}^{4}$ are the restriction of a global order $d$ rotation on $\tilde{B}^{5}$ fixing a properly embedded unknotted 3-ball. If $(\tilde{S},\rho)$ and $(\tilde{S}^{\prime},\rho^{\prime})$ are directed, we say that a concordance between them is directed if the orientations on $\operatorname{Fix}(\rho)$ and $\operatorname{Fix}(\rho^{\prime})$ agree. Directed equivariant concordance is then an equivalence relation on the set of directed periodic 2-knots.
@@
Let $C_{2}^{d}$ be the set of directed $d$-periodic 2-knots. The equivariant concordance group $\mathcal{C}_{2}^{d}$ of $d$-periodic 2-knots is the quotient of the set $C_{2}^{d}$ by directed equivariant concordance, with group multiplication given by equivariant connected sum.